1. 4(3a^2)^2
2. A^10/A^10

simplify

Answer by stanbon(75887)

1. 4(3a^2)^2 = 4(9a^4) = 36a^4
2. A^10/A^10 = A^0 = 1
=====================================
Cheers,
Stan H.